# InvisibleInk: High-Utility and Low-Cost Text Generation with Differential Privacy

### v1

@@ 4 InvisibleInk: DP Text Generation under Strict Compute Budgets
DClip: A Targeted Clipping Function. The private next-token logits ${\bm{\phi}}_{i}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})$ encode not just the sensitive information in the reference ${\bm{r}}_{i}$, but also general information about language, grammar and syntax, semantics, a degree of world knowledge and common sense, and style/tone of the preceding tokens [e.g. 65, 66, 67]. Clipping the next-token logits involves paying a privacy cost for this general non-private information, leading to wasted privacy budget.
@@
where $C>0$ is a specified clip norm, and $\mathsf{clip}_{C}({\bm{\phi}})$ projects each coordinate of ${\bm{\phi}}$ onto the interval $[-C,C]$. The clipped and aggregated logits $\bar{\bm{\phi}}$ can then be a drop-in replacement in Eq. 2 is
@@
We observe that the spread of ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ is typically much smaller than that of ${\bm{\phi}}_{i}$, as shown in Fig. 3. This is because the non-private logits ${\bm{\phi}}_{\mathsf{pub}}$ already contain a strong prior on language. This lets us utilize a much smaller clip norm $C$ without distorting the model outputs, and hence a smaller compute cost $B$ to maintain a temperature of $\tau\propto C/B$ close to 1.
@@
1:LLM logit API $\phi(\cdot\,|\,\cdot)$, vocabulary $V$, query ${\bm{q}}$, sensitive references ${\bm{R}}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{B}\}$, max text length $T$, clip norm $C$, temperature $\tau$, top-$k$ parameter, initial generation ${\bm{x}}=\emptyset$
[…]
3:    For $i\in[B]$, set ${\bm{\phi}}_{i}=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$ and ${\bm{\phi}}_{\mathsf{pub}}=\phi(\,\cdot\,|\,{\bm{q}},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$triangleright LLM calls
4:    $V_{k}^{+}=\textsc{ExpandedTopVocabulary}\big{(}{\bm{\phi}}_{\mathsf{pub}},k,C,B\big{)}$triangleright Vocabulary for Top-$k+$ sampling
5:    Set $\bar{{\bm{\phi}}}={\bm{\phi}}_{\mathsf{pub}}+\frac{1}{B}\sum_{i=1}^{B}\;\mathsf{clip}_{C}({\bm{\phi}}_{i},{\bm{\phi}}_{\mathsf{pub}})$triangleright Aggregated clipped logits
6:    Sample $x_{t}\;\sim\;\mathsf{softmax}(\bar{\bm{\phi}}[V_{k}^{+}]/\tau)$triangleright Exponential mechanism over $V_{k}^{+}$
[…]
@@
9:    Return $\{y\in V\,:\,{\bm{\phi}}_{\mathsf{pub}}(y)\geq\ell-2C/B\}$, where $\ell$ is the $k$th largest entry of ${\bm{\phi}}_{\mathsf{pub}}$
@@
Vanilla (non-private) top-$k$ sampling from one logit vector ${\bm{\phi}}_{i}=\phi_{i}(\,\cdot\,|{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})$ would restrict the vocabulary to $V_{k}({\bm{\phi}}_{i})$. However, in our case, we have $B$ such vectors ${\bm{\phi}}_{1},\ldots,{\bm{\phi}}_{B}$ which have to be combined to get the next token. Our goal is to extend the truncation to union of the top-$k$ contributions of these logit vectors in Eq. 4. This is the set $\bar{V}_{k}:=\cup_{i=1}^{B}V_{k}({\bm{\phi}}^{\mathsf{clip}}_{i})$, where we obtain the standalone contribution ${\bm{\phi}}^{\mathsf{clip}}_{i}$ of ${\bm{\phi}}_{i}$ in Eq. 4 by assuming ${\bm{\phi}}_{j}={\bm{\phi}}_{\mathsf{pub}}$ contains no private information for $j\neq i$:
@@
Unfortunately, restricting our vocabulary to this privacy-sensitive top-$k$ set $\bar{V}_{k}$ fails to satisfy DP. For example, if moving from a set ${\bm{R}}$ of sensitive references to an adjacent ${\bm{R}}^{\prime}$ causes a token to jump into $\bar{V}_{k}$ from outside (or vice-versa), we get a privacy loss of +$\infty$+ ∞. An alternative that has been explored in previous work is to simply sample from the top-$k$ tokens from the public logits ${\bm{\phi}}_{\mathsf{pub}}$. There is no privacy cost here as this decision is independent of the private logits. However, this approach may fail to generate tokens that are rare in the public (pretraining) data but common in the sensitive data ${\bm{R}}$.
Instead, our approach is to build a tight superset of $\bar{V}_{k}$ based purely on the public logits ${\bm{\phi}}_{\mathsf{pub}}$. Let $\ell$ be the top-$k$ threshold of the public logits ${\bm{\phi}}_{\mathsf{pub}}$. Then, all possible top-$k$ clipped logits of ${\bm{\phi}}^{\mathsf{clip}}_{i}$ are bounded from below by $\ell-C/B$. Thus, the top-$k$ indices of ${\bm{\phi}}^{\mathsf{clip}}_{i}$ for each $i\in[B]$ are guaranteed to lie within the set $V_{k}^{+}:=\{y\in V\,:\,{\bm{\phi}}_{\mathsf{pub}}(y)\geq\ell-2C/B\}$. We refer to sampling from this expanded vocabulary set $V_{k}^{+}$ as Top-$k+$ sampling.
We perform the truncated sampling only over this expanded top-$k$ set $V_{k}^{+}$. Note that this set is constructed without using any information about the private logits ${\bm{\phi}}_{1},\ldots,{\bm{\phi}}_{B}$, so it can be used freely without any privacy cost. We find in §5 that the top-$k$ expansion adds fewer than 10 tokens in practical settings, making it a very tight superset of the private top-$k$ vocabulary $\bar{V}_{k}$.
InvisibleInk: Text Generation & Privacy Accounting. Our approach wraps the above two components in an exponential mechanism-based outer loop, as given in Algorithm 1. Notice crucially that Algorithm 1 does not involve any data-dependent privacy mechanisms. This lets us give a direct zCDP bound, building directly upon the zCDP analysis of the exponential mechanism [63].
@@
Algorithm 1 with a maximum token budget $T$, a clipping threshold $C$, a set ${\bm{R}}$ of $B=|{\bm{R}}|$ references, and temperature $\tau$ satisfies $\rho_{\mathsf{seq}}$-zCDP with $\rho_{\mathsf{seq}}={TC^{2}}/{(2B^{2}\tau^{2})}$.
When running the algorithm in practice, we assume that we are given the privacy budget $\rho_{\mathsf{seq}}$, the maximum sequence length $T$, and the number of references $B$ (which fixes the compute budget). We recommend a default temperature of $\tau\approx 1$. Then, Theorem 2 lets us set the clip norm $C=B\tau\sqrt{2\rho_{\mathsf{seq}}/T}$ to get the desired privacy guarantee (independent of top-$k$ parameter).
Another advantage of our approach is that the generations gracefully fall back on that of the public model if the computation budget (i.e., the size $B$) is too small for a given privacy budget $\rho_{\mathsf{seq}}$. Indeed, in this case, we get a very small clip norm, so that $\bar{\bm{\phi}}_{i}\approx{\bm{\phi}}_{\mathsf{pub}}$.
This nice property is not satisfied by existing baselines. For example, as the privacy budget or the computation budget gets small, Amin et al. [12] requires a small clip norm $C$, reducing their method to uniform sampling over the full vocabulary (see Fig. 3). This leads to gibberish text, as the vocabulary contains many word pieces rather than full words.777 For example, the word “differential” might be made up of two tokens “differ” and “-ential”. Sampling uniformly at random could yield the “-ential” token without an appropriate preceding token, producing gibberish. Similarly, AdaPMixED [18] has a data-dependent privacy guarantee that depends on the number of tokens sampled per generation. As the privacy budget or the computation budget gets smaller, the length of synthetic generations falls very sharply. In contrast, by falling back on the public models, our method always returns coherent text (although it might be somewhat generic when $B$ is very small, as it does not use the sensitive references).
Finally, we note that it might be possible to improve the dependence of Theorem 2 on the token budget $T$ with regularization [e.g. 68, 69]; we leave this for future work.
@@ 5 Experiments
We evaluate InvisibleInk on DP synthetic text generation. We will release our code upon publication.
5.1 Experimental Setup
Given a reference dataset $\mathcal{D}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{N}\}$, our task is to create $n$ synthetic texts ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ that are statistically similar to $\mathcal{D}$ while satisfying a given $\rho_{\mathsf{seq}}$-zCDP guarantee with respect to $\mathcal{D}$. We generate synthetic example ${\bm{x}}_{j}$ based on a batch ${\bm{R}}_{j}=({\bm{r}}_{jB+1},\ldots,{\bm{r}}_{(j+1)B})$ of sensitive references with a query ${\bm{q}}$ instructing a generation similar to the given references. By parallel composition, each example ${\bm{x}}_{j}$ must also satisfy $\rho_{\mathsf{seq}}$-zCDP. Note that the batch size $B$ fixes the computation budget: we allow $(B+1)$ LLM next-token inference calls per generated token. The batching limits us to $n\leq\lfloor N/B\rfloor$ synthetic examples. See §D for extensive details on the task and experimental setup; we summarize the key details here.
Dataset Domain Text Type Size Avg. Length # Generations Max. Length MIMIC-IV Medical Discharge Summaries 311K 298.4 1000 500 Yelp Commercial Reviews 1.9M 145.2 500 200 TAB Legal Court Case Texts 1013 387.7 100 500
@@ 5.1 Experimental Setup
Datasets. We experiment with three datasets in clinical, legal, and commercial domains, see Tab. 1. MIMIC-IV [70] is a de-identified collection of medical text associated with ICU patients; we use the discharge summaries, which contain sensitive diagnosis, treatment, and medication information. Further, this dataset is legally unavailable for LLM pretraining, making it valuable in modeling real-world distribution-shift sensitive domains. The Text Anonymization Benchmark (TAB) [71] contains $N=1013$ (training set) court case notes (in English) from the European Court of Human Rights (ECHR) with personal identifiers and other confidential attributes. This low-resource dataset only admits extremely small batch sizes $B$ (e.g., $B=127$ gives us a maximum of $8$ synthetic texts). The Yelp Reviews dataset [72] contains user-generated reviews and ratings for businesses with personal opinions and location references, and is a standard benchmark in DP text generation [12, 36].
@@
Baselines and Hyperparameters. We compare InvisibleInk (with and without Top-$k+$ sampling) with the prior white-box SoTA methods Amin et al. [12] for DP text generation and AdaPMixED [18] for DP next-token prediction. We iteratively sample from the predicted next-token distribution of AdaPMixED to adapt it to DP text generation. For all methods, we fix the batch size $B$ as it fixes the computation budget. We select other hyperparameters, such as the clip norm $C$ for InvisibleInk and Amin et al. [12] to achieve a desired privacy guarantee. We sweep the temperature over a grid for Amin et al. [12] and report values for the best temperature. Similarly, we sweep the top-$k$ parameter over a grid for InvisibleInk (with temperature fixed) and report values for the best parameter.
The method of Amin et al. [12] has a sensitivity of $C/B$ under the zero-out adjacency, compared to $2C/B$ under the replace-by-null adjacency (see Property 9 of §B). Yet, we give it an advantage by (incorrectly) taking the sensitivity to be $C/B$ under the replace-by-null adjacency. AdaPMixED originally gives a data-dependent DP guarantee under the add-or-remove adjacency in [18]. While AdaPMixED can be directly modified to use replace-by-null adjacency, we find that the utility trends under both adjacency notions are qualitatively similar: we report the utility for add-or-remove adjacency here and compare it with replace-by-null adjacency in §B.3.
@@
Metrics. We evaluate the statistical similarity between the generated text and reference distribution with MAUVE scores [16, 76, 17] and the gap between the perplexities of generated and reference texts (denoted $\Delta$PPL). The MIMIC discharge notes contain medical named entities such as names of medical conditions, procedures, and medicines. We plot the number of such entities identified by the MedNER model [77]. Finally, we also plot the average generation length, as some methods can fail to produce sufficient tokens before exhausting the privacy budgets.
Figure 4: Utility-compute tradeoffs at $(\varepsilon=10,\delta=10^{-6})$ DP on each dataset across varying compute budget from $B\in\{2,\ldots,128\}$. InvisibleInk can produce text that matches or exceeds the baselines at a fraction of the compute. The baselines do not even work for the low-resource TAB dataset at a small batch size $B=8$.

### v2

@@ 4 InvisibleInk: DP Text Generation under Strict Compute Budgets
DClip: A Targeted Clipping Function. The private next-token logits ${\bm{\phi}}_{i}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})$ encode not just the sensitive information in the reference ${\bm{r}}_{i}$, but also general information about language, grammar and syntax, semantics, a degree of world knowledge and common sense, and style/tone of the preceding tokens [e.g. 73, 74, 75]. Clipping the next-token logits involves paying a privacy cost for this general non-private information, leading to wasted privacy budget.
@@
where $C>0$ is a specified clip norm, and $\mathsf{clip}_{C}({\bm{\phi}})$ projects each coordinate of $\bm{\phi}$} onto the interval [C,C][-C,C]. The clipped and aggregated logits $\bar{\bm{\phi}}$ can then be a drop-in replacement in Eq.˜2 is
@@
Fig.˜3 shows that the spread of ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ is typically much smaller than that of ${\bm{\phi}}_{i}$, since ${\bm{\phi}}_{\mathsf{pub}}$ already contains a strong prior on language. We utilize a smaller clip norm $C$ without distorting the model outputs, and hence a smaller compute cost $B$ to maintain a temperature of $\tau\propto C/B$ close to 1.
@@
1:LLM logit API $\phi(\cdot\,|\,\cdot)$, vocabulary $V$, query $\bm{q}$}, sensitive references ${\bm{R}}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{B}\}$, max text length $T$, clip norm $C$, temperature $\tau$, top-$k$ parameter, initial generation $\bm{x}$$\emptyset$
2:for $t\;=\;1,\ldots,T$ do
3:  For $i\in[B]$, set ${\bm{\phi}}_{i}=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$ and ${\bm{\phi}}_{\mathsf{pub}}=\phi(\,\cdot\,|\,{\bm{q}},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$ $\triangleright$ LLM calls
4:  $V_{k}^{+}=\textsc{ExpandedTopVocabulary}\big({\bm{\phi}}_{\mathsf{pub}},k,C,B\big)$ $\triangleright$ Vocabulary for Top-$k+$ sampling
5:  Set $\bar{{\bm{\phi}}}={\bm{\phi}}_{\mathsf{pub}}+\frac{1}{B}\sum_{i=1}^{B}\;\mathsf{clip}_{C}({\bm{\phi}}_{i},{\bm{\phi}}_{\mathsf{pub}})$ $\triangleright$ Aggregated clipped logits
6:  Sample $x_{t}\;\sim\;\mathsf{softmax}(\bar{\bm{\phi}}[V_{k}^{+}]/\tau)$ $\triangleright$ Exponential mechanism over $V_{k}^{+}$
7:  Yield next token $x_{t}$, append it to $\bm{x}$}, and break if $x_{t}=\texttt{<eos>}$ $\triangleright$ Generated Token
@@
9:  Set $\ell$ to be the $k$th largest entry of ${\bm{\phi}}_{\mathsf{pub}}$ $\triangleright$ top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$
10:  Return $\{y\in V\,:\,{\bm{\phi}}_{\mathsf{pub}}(y)\geq\ell-2C/B\}$ $\triangleright$ Expand top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$ by $2C/B$
@@
Vanilla (non-private) top-$k$ sampling from one logit vector ${\bm{\phi}}_{i}=\phi_{i}(\,\cdot\,|{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})$ would restrict the vocabulary to $V_{k}({\bm{\phi}}_{i})$. However, in our case, we have $B$ such vectors ${\bm{\phi}}_{1},\ldots,{\bm{\phi}}_{B}$ which have to be combined to get the next token. Our goal is to extend the truncation to union of the top-$k$ contributions of these logit vectors in Eq.˜4.
To get the contribution of the logit vector ${\bm{\phi}}_{i}$ to Eq.˜4, we plug in ${\bm{\phi}}_{j}={\bm{\phi}}_{\mathsf{pub}}$ for $j\neq i$, since ${\bm{\phi}}_{j}$ does not contain any private information about ${\bm{\phi}}_{i}$. Thus, we define its contribution as
@@
Then, the top-$k$ vocabulary for private generation is simply the union $\bar{V}_{k}:=\cup_{i=1}^{B}V_{k}({\bm{\phi}}^{\mathsf{clip}}_{i})$.
Unfortunately, restricting our vocabulary to this privacy-sensitive top-$k$ set $\bar{V}_{k}$ fails to satisfy DP. Changing the dataset from $\bm{R}$} to an adjacent $\bm{R}$$^{\prime}$, can cause a token $y\in V$ to “jump out” of the top-$k$ set $\bar{V}_{k}$. Its probability of being generated then goes from a non-zero value to a zero value, leading to a privacy loss of $\log(Q(y|{\bm{q}},{\bm{R}},{\bm{x}}_{<t})/Q(y|{\bm{q}},{\bm{R}}^{\prime},{\bm{x}}_{<t}))=\infty$ (see [e.g. 70, Def. 2] for definitions).
An alternative that has been explored in previous work is to simply sample from the top-$k$ tokens from the public logits ${\bm{\phi}}_{\mathsf{pub}}$. This incurs no privacy cost as the restricted vocabulary is independent of the private logits. However, this approach may fail to generate tokens that are rare in the public (pretraining) data but common in the sensitive data $\bm{R}$}. Such tokens are likely to be relevant to the data generation domain, and it is thus desirable to recover them.
Instead, we build a tight superset of $\bar{V}_{k}=\cup_{i=1}^{B}V_{k}({\bm{\phi}}_{i}^{\mathsf{clip}})$ based purely on the public logits ${\bm{\phi}}_{\mathsf{pub}}$. Observe that clipping ensures $|\phi_{i}^{\mathsf{clip}}(y)-\phi_{\mathsf{pub}}(y)|\leq C/B$ for each $i$. So, each $y\in V_{k}({\bm{\phi}}_{\mathsf{pub}})$ satisfies ${\bm{\phi}}_{i}^{\mathsf{clip}}(y)+C/B\geq\ell$, where $\ell$ is the top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$. Thus, the top-$k$ threshold $\ell_{i}$ of ${\bm{\phi}}_{i}^{\mathsf{clip}}$ is bounded as $\ell_{i}\geq\ell-C/B$. It thus follows that:
$y\in V_{k}({\bm{\phi}}_{i}^{\mathsf{clip}})\iff{\bm{\phi}}_{i}^{\mathsf{clip}}(y)\geq\ell_{i}\implies{\bm{\phi}}_{\mathsf{pub}}(y)+\tfrac{C}{B}\geq\ell-\tfrac{C}{B}\,.$ (5)
In other words, the private top-$k$ vocabulary $V_{k}({\bm{\phi}}_{i}^{\mathsf{clip}})$ always lies in the set $V_{k}^{+}:=\{y\in V\,:\,{\bm{\phi}}_{\mathsf{pub}}(y)\geq\ell-2C/B\}$. Since this is true for each $i\in[B]$, we have that $V_{k}^{+}$ is a superset of $\bar{V}_{k}=\cup_{i=1}^{B}V_{k}({\bm{\phi}}_{i}^{\mathsf{clip}})$. We refer to sampling from the set $V_{k}^{+}$ as Top-$k+$ sampling.
Since $V_{k}^{+}$ is constructed without using the private logits, sampling from it does not incur any privacy cost. We find in §5 that $\left|V_{k}^{+}-V_{k}({\bm{\phi}}_{\mathsf{pub}})\right|\approx 10\ll|V|$, i.e. $V_{k}^{+}$ is indeed a tight superset of $\bar{V}_{k}$. We further develop the intuition on Top-$k+$ sampling using a toy example in §C.3.
InvisibleInk: Text Generation & Privacy Accounting. Our approach wraps the above two components in an exponential mechanism-based outer loop; see Algorithm˜1. Crucially, Algorithm˜1 does not involve any data-dependent privacy mechanisms. This lets us give a zCDP bound, building directly upon the zCDP analysis of the exponential mechanism [71] and adaptively composing [70] the per-token zCDP guarantee over $T$ tokens. Adaptive composition ensures that previously generated tokens can be reused safely in future steps without additional privacy cost; see §C for proof.
@@
Algorithm˜1 with a maximum token budget $T$, a clipping threshold $C$, a set $\bm{R}$} of $B=|{\bm{R}}|$ references, and temperature $\tau$ satisfies $\rho_{\mathsf{seq}}$-zCDP with $\rho_{\mathsf{seq}}={TC^{2}}/{(2B^{2}\tau^{2})}$.
When running the algorithm in practice, we assume that we are given the privacy budget $\rho_{\mathsf{seq}}$, the maximum sequence length $T$, and the number of references $B$ (which fixes the compute budget). We recommend a default temperature of $\tau\approx 1$. Then, Theorem˜2 lets us set the clip norm $C=B\tau\sqrt{2\rho_{\mathsf{seq}}/T}$ to get the desired privacy guarantee (independent of the top-$k$ parameter).
Another advantage of our method is that generation gracefully falls back on that of the public model if the compute or privacy budget is too small. In this case, we get a very small $C$ and $\bar{\bm{\phi}}_{i}\approx{\bm{\phi}}_{\mathsf{pub}}$. This nice property is not satisfied by existing baselines. For example, as the privacy or computation budget gets smaller, [13] requires a small clip norm $C$, reducing their method to uniform sampling over the full vocabulary (see Fig.˜3). This leads to gibberish text. [^6] For example, the word “differential” might be made up of two tokens “differ” and “-ential”. Sampling uniformly at random could yield the “-ential” token without an appropriate preceding token, producing gibberish. Similarly, AdaPMixED [19] has a data-dependent privacy guarantee that depends on the number of tokens sampled per generation. As the privacy or computation budget gets smaller, the length of synthetic generations falls sharply.
Finally, we note that it might be possible to improve the dependence of Theorem˜2 on the token budget $T$ with regularization [e.g. 76, 77]; we leave this for future work.
@@ 5 Experiments
Setup. Given a reference dataset $\mathcal{D}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{N}\}$, our task is to create $n$ synthetic texts ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ that are statistically similar to $\mathcal{D}$ while satisfying a given $\rho_{\mathsf{seq}}$-zCDP guarantee with respect to $\mathcal{D}$. We generate synthetic example ${\bm{x}}_{j}$ based on a batch ${\bm{R}}_{j}=({\bm{r}}_{jB+1},\ldots,{\bm{r}}_{(j+1)B})$ of sensitive references with a query $\bm{q}$} instructing a generation similar to the given references. By parallel composition, each example ${\bm{x}}_{j}$ must also satisfy $\rho_{\mathsf{seq}}$-zCDP. Since the batch size $B$ fixes the computation budget, we allow $(B+1)$ LLM next-token inference calls per generated token. This also limits us to $n\leq\lfloor N/B\rfloor$ synthetic examples. See §D for additional details on the task and experimental setup.
We report a central DP guarantee for a single hyperparameter setting to generate $n$ synthetic text samples; the randomized mechanism here is a set of $n$ autoregressive text generations from LLMs, which output a sequence of tokens [^7] The token string ends in an <EOS> token unless the number of tokens generated reaches the limit $T$ first. each. We propose heuristics for choosing optimal hyperparameters in §5.3 in lieu of accounting for privacy lost in hyperparameter tuning. Throughout this work, we report example-level (an example is a single text sample drawn from the reference dataset) $\rho$-zCDP guarantees (converted to $(\varepsilon,\delta)$-DP where necessary) under the replace-by-null adjacency; see §B for details. We release our code at: https://github.com/cerai-iitm/InvisibleInk-Experiments.
Datasets. We experiment with three datasets in clinical, legal, and commercial domains, see Tab.˜1. MIMIC-IV-Note [78, 79] is a de-identified collection of medical text associated with ICU patients; we use the discharge summaries, which contain sensitive diagnosis, treatment, and medication information. Further, this dataset is legally unavailable for LLM pretraining, making it valuable in modeling real-world distribution-shift sensitive domains. The Text Anonymization Benchmark (TAB) [80] contains $N=1013$ (training set) court case notes (in English) from the European Court of Human Rights (ECHR) with personal identifiers and other confidential attributes. This low-resource dataset only admits extremely small batch sizes $B$ (e.g., $B=127$ gives us a maximum of $8$ synthetic texts). The Yelp Reviews dataset [81] contains user-generated reviews and ratings for businesses with personal opinions and location references, and is a standard benchmark in DP text generation [13, 41]. We intentionally choose real privacy-sensitive datasets (e.g. MIMIC, TAB) datasets with long-form text, rather than standard internet datasets (AGNews, IMDB, etc.) that are well-represented in the pretraining data of most LLMs [82, 83]. See §D.2 for further discussion.
Dataset Domain Text Type Size Avg. Length # Generations Max. Length MIMIC-IV-Note Medical Discharge Summaries 311K 298.4 1000 500 Yelp Commercial Reviews 1.9M 145.2 500 200 TAB Legal Court Case Texts 1013 387.7 100 500
@@
Baselines and Hyperparameters. We compare InvisibleInk (with and without Top-$k+$ sampling) with the prior white-box SoTA methods Amin et al. [13] for DP text generation and AdaPMixED [19] for DP next-token prediction. We iteratively sample from the predicted next-token distribution of AdaPMixED to adapt it to DP text generation. For all methods, we fix the batch size $B$ as it fixes the computation budget. We select other hyperparameters, such as the clip norm $C$ for InvisibleInk and Amin et al. [13] to achieve a desired privacy guarantee. See §D.5 for further details.
The method of Amin et al. [13] has a sensitivity of $C/B$ under the zero-out adjacency, compared to $2C/B$ under the replace-by-null adjacency (see Property˜9 of §B). Yet, we give it an advantage by (incorrectly) taking the sensitivity to be $C/B$ under the replace-by-null adjacency. AdaPMixED originally gives a data-dependent DP guarantee under the add-or-remove adjacency in [19]. Since we find that the utility trends under both adjacency notions are qualitatively similar (see §E.5), we modify AdaPMixED directly and report the results for replace-by-null adjacency here.
@@
In addition to the baselines discussed above, we also compare InvisibleInk with the API-access method of Augmented Private Evolution (AugPE) [41]. This more restrictive setting naturally makes privatization harder. AugPE makes two types of LLM inference calls: generation and paraphrasing. We cannot directly compare the computational cost of InvisibleInk to AugPE in an apples-to-apples comparison based on the number of next-token inference calls per generated token, since it is calibrated using the number of paraphrasing iterations made per generated sequence. Instead, we directly use the wall-clock runtime of both methods as a measure of computation cost. We also allow AugPE an advantage of $\sqrt{2}$ in the sensitivity under replace-by-null adjacency (see §B.3.3). For additional results and detailed generation settings, see §E.1.
[…]
Figure 4: Utility-compute tradeoffs at $(\varepsilon=10,\delta=10^{-6})$ DP on each dataset across varying compute budget from $B\in\{1,3,\ldots,127\}$. Results reported over 33 runs with $95\%$ confidence intervals (see §D.6) for MIMIC and 11 run for Yelp/TAB datasets. InvisibleInk can produce text that matches or exceeds the baselines at a fraction of the compute. The baselines do not even work for the low-resource TAB dataset at a small batch size $B=7$.